OpenStudy (anonymous):

can someone help me with this(3/4)^-1+(2/3)

OpenStudy (anonymous):

Do I distribute the -1 into the fraction?

OpenStudy (anonymous):

how do I view the reply?

OpenStudy (anonymous):

I would bvery grateful if some one can help me become fimilar with this site. It seem like it can help a ton of people. I was wondering if someone could tell me how the view the replies?

OpenStudy (anonymous):

Hello! So whenever you have something to the negative one power, like what you have, this is (in your case) indicating that you can "flip" whatever is on the numerator to the denominator or vice versa. So technically with your problem you would first distribute the negative one like so: (3^-1)/(4^-1)+2/3 So then using what I said previously, it will now look like this 4/3+2/3 I leave the rest up to you :D.

OpenStudy (anonymous):

so I have \[4/3+2/3=6/3=2\]

OpenStudy (anonymous):

Exactly. So did you understand how you would do the process if you got another problem similar to this one, say (2^-1)/5+3/10?

OpenStudy (anonymous):

No how did you get the 4/3

OpenStudy (anonymous):

Okay so you know how I said you could distribute the -1, let's start from there. So we have (3^-1)/(4^-1). Now the 3 in the numerator, since it is to the power of the negative one, can now be flipped to the denominator. So we have: 1/(3*(4)^-1). So we took care of the three, because once you flip it, the negative sign can go away. Now time to take care of the 4^-1. Since it is in the denominator and 4 is to the power of negative one, we can flip the four to the numerator (but be sure to leave the three, that is no longer to a negative power, alone). So now we have 4/3. Did that help a little more?

OpenStudy (anonymous):

Yes a ton thank you

OpenStudy (anonymous):

So for the example that I typed what would you think you should do first?

OpenStudy (anonymous):

^mistype: what do you think you should do first**

OpenStudy (anonymous):

would I go (2^-1)/(4^-1)/5+3?10 ??? Then

OpenStudy (anonymous):

???

OpenStudy (anonymous):

So here is the original: (2^-1)/5+3/10 I would first look at how only the negative two is to the negative power.

OpenStudy (anonymous):

So that is the only one that will be changing locations, so to speak. :D

OpenStudy (anonymous):

so I would have (2^-1)(3^-1)+3/10

OpenStudy (anonymous):

will that give me 3/5 + 3/10?

OpenStudy (anonymous):

I get it hold on

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

5/2 + 3/5+ = 8/7 is this right

OpenStudy (anonymous):

(2^-1)/5+3/10 So I would only flip the 2 to the denominator. So since multiplication is the opposite of division we will have 1 in the numerator of the fraction and 2*5 on the denominator in the first fraction. In otherwords, 1/(2*5)+3/10

OpenStudy (anonymous):

so i will have 1/10 + 3/10 =3/10

OpenStudy (anonymous):

1+3=?

OpenStudy (anonymous):

sorry Im adding 4

OpenStudy (anonymous):

4/20

OpenStudy (anonymous):

which is 5

OpenStudy (anonymous):

late night sorry. I really appricate you haelping me out.

OpenStudy (anonymous):

late night sorry. I really appricate you haelping me out.

OpenStudy (anonymous):

Close again. It's actually 4/10 (you only add stuff in the numerator if the denominator is the same number).

OpenStudy (anonymous):

No Problem, by the way.

OpenStudy (anonymous):

that is true I forgot about that

OpenStudy (anonymous):

So key things to keep in mind. Numbers where some number in the numerator is to the -1 power like 4^-1 will equal 1/4. Numbers like 1/(6^-1) where the number is in the denominator will equal 6.

OpenStudy (anonymous):

thanks for the tip

OpenStudy (anonymous):

Mhm. I hope that helped you a little.